Prove the foreign investment multiplier is equal to 1 all over mpm + mps (1/mpm+mps)


1
Expert's answer
2021-06-01T11:28:32-0400

We derive the foreign investment multiplier as follows:

The national income identity in an open economy is given by "Y=C+I+C-M"

Y is the national income, C national consumption,I total investment, X exports and M thee imports.

This relationship can also be arranged as

"Y-C=I+C-M \\ or\\S=I+C-M(S=Y-C)\\S+M=I+X"

S denotes savings

At equilibrium level of income "S+M=I+X"

In open economy

"I=S\\ I_d+I_f=S....(1)"

Id denotes domestic investment while If denotes foreign investment.

"I_f=X-M....(2)"

Substitute (2) into( 1)

"I_d+X-M-S \\ or\\ I_d+X=S+M"

Which is the equilibrium condition of national income in open economy.

The foreign trade investment multiplier coefficient (Kf)is equal to:

"K_f=\frac{∆Y}{∆X}\\ and\\∆X=∆S+∆M"

Divide both sides by ∆Y

"\frac{∆X}{∆Y}=\frac{∆S+∆M}{∆Y}"


"\frac{∆Y}{∆X}=\frac{∆Y}{∆S+∆M}"

"K_f=\frac{∆Y}{∆S+∆M}"

"K_f=\frac{1}{\frac{∆S}{∆Y}+\frac{∆M}{∆S}}"

"K_f=\frac{1}{MPS+MPM}"